OpenStudy (anonymous):

if two zeroes of the poly.x^4-7x*3+17x^2-17x+6 are 3 and 1.find the other zeroes of the poly.

OpenStudy (anonymous):

@apoorvk

OpenStudy (anonymous):

@lgbasallote

OpenStudy (lgbasallote):

did you know that since 3 and 1 are zero then: x = 3 and x = 1 therefore (x-3) and (x-1) are factors of the polynomial

OpenStudy (anonymous):

k dennnn

OpenStudy (lgbasallote):

so (x-3)(x-1) = x^2 - x - 3x + 3 = x^2 - 4x = 3 now divide the polynomial given with this

OpenStudy (anonymous):

frm where u gt x^2 - x - 3x + 3

OpenStudy (apoorvk):

If the zeroes of this polynomial are 3 and 1, that means (x-3) and (x-1) are factors of this polynomial. Divide the polynomial by any of the factors, and obtain a quotient in terms of 'x'. This quotient will be a cubic. Then divide this quotient again, this time by the other factor. You 'll get a quadratic. Factorise this quadratic and get your roots!

OpenStudy (lgbasallote):

FOIL (x-3)(x-1)

OpenStudy (apoorvk):

or as @lgbasallote says, divide it by (x-1)(x-3) just once, and get the quadratic.

OpenStudy (anonymous):

so x=3 and x=1

OpenStudy (anonymous):

????????????????????????/

OpenStudy (anonymous):

ma bulb is fused it is nt glowing wht to do

OpenStudy (lgbasallote):

where did you get lost??

OpenStudy (anonymous):

c i gt (x-3) and (x-1) den wht did u do? @lgbasallote

OpenStudy (lgbasallote):

(x-3) and (x-1) are factors of the polynomial in the question..do you get that?

OpenStudy (anonymous):

yssss after st its gne

OpenStudy (anonymous):

step by step can u shw

OpenStudy (lgbasallote):

next...you multiply them... (x-3) * (x-1)

OpenStudy (anonymous):

so we will gt x*-9??????

OpenStudy (lgbasallote):

no..remember FOIL \[(x * x) + (x * -1) + (-3 * x) + (-3 * -1)\] got it?

OpenStudy (anonymous):

* means????

OpenStudy (anonymous):

multiply?

OpenStudy (anonymous):

is it @lgbasallote

OpenStudy (lgbasallote):

yes

OpenStudy (anonymous):

so we will gt x^2-x-3x+3 is it????

OpenStudy (lgbasallote):

yup..then combine -x and -3x

OpenStudy (apoorvk):

Great work, lgba! :)

OpenStudy (anonymous):

dennn plzz shw it den only i will gt it

OpenStudy (anonymous):

a big thx to @lgbasallote

OpenStudy (lgbasallote):

x^2 - x - 3x + 3 = x^2 - 4x + 3

OpenStudy (lgbasallote):

-x - 3x is like -1 - 3...just with an x

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

den

OpenStudy (lgbasallote):

divide... \[\large (x^4 - 7x^3 + 17x^2 - 17x + 6) \div (x^2 - 4x + 3)\]

OpenStudy (anonymous):

hence the another roots are 1AND 2

OpenStudy (anonymous):

k thx to @lgbasallote

OpenStudy (anonymous):

so wht will we gt i hav to confirm dts y @lgbasallote

OpenStudy (lgbasallote):

from the division?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

i gt 0 as d remainder and the g(x)=x^2-3x+2

OpenStudy (anonymous):

correct? finished d problem

OpenStudy (anonymous):

@lgbasallote

OpenStudy (anonymous):

is it correct @lgbasallote

OpenStudy (lgbasallote):

wait im checking lol

OpenStudy (anonymous):

k

OpenStudy (anonymous):

u r awsome dude mainly i depend on @apoorvk bt 2day i gt new frnd u @lgbasallote thxxxxxxxxxxx

OpenStudy (lgbasallote):

yup it's right

OpenStudy (lgbasallote):

so factor out \(x^2 - 3x + 2\)

OpenStudy (anonymous):

ohhhh god u do dis :P

OpenStudy (lgbasallote):

\[x^2 - 3x + 2 = (x-2)(x-1)\] lol you need to learn factoring ;)

OpenStudy (lgbasallote):

now do you know how to find the roots?

OpenStudy (apoorvk):

@Parvathysubhash - @lgbasallote is miles ahead and tonnes better than me - mind it ;)

OpenStudy (anonymous):

it will b 2 and 1 and we hav to find d sum and d product rhyt

OpenStudy (lgbasallote):

lol @apoorvk :P

OpenStudy (anonymous):

anyway u bth r awsome @apoorvk and @lgbasallote

OpenStudy (lgbasallote):

it's just looking for the other zeroes so x = 2 and x = 1 are the last two zeroes

OpenStudy (lgbasallote):

got it?

OpenStudy (anonymous):

so nw i hav to find sum and product rhyt

OpenStudy (lgbasallote):

why??

OpenStudy (anonymous):

sum of the zeroes and product of the zeroes we hav to na? no need?

OpenStudy (lgbasallote):

isn't the question just asking what are the other two zeroes? is there still another question looking for sum and product??

OpenStudy (anonymous):

kkkkk

OpenStudy (lgbasallote):

now...i sleep lol

OpenStudy (anonymous):

k

OpenStudy (anonymous):

gudni8 if u wnt u stay i'll ask u many question and u answer dt :P

OpenStudy (apoorvk):

We stay, you ask :P

OpenStudy (anonymous):

thxxxxx